### v2

# Thomason cohomology and Quillen’s Theorem A

Abstract.
Given a functor $\varphi:\mathcal{C}\to\mathcal{D}$ between two small categories, there is a homotopy equivalence $\kappa:\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\to N\mathcal{C}$ where $N(\varphi/-)$ is the functor which sends every object $d$ in $\mathcal{D}$ to the nerve of the comma category $\varphi/d$. We prove that the homotopy equivalence $\kappa$ induces an isomorphism on cohomology with coefficients in any coefficient system. As a consequence, we obtain a version of Quillen’s Theorem A for the Thomason cohomology of categories. We also construct a spectral sequence for the Thomason cohomology of the Grothendieck construction $\int_{\mathcal{D}}F$ of a functor $F:\mathcal{D}\to Cat$ using the isomorphism in the main theorem.
Key words and phrases:
Cohomology of small categories, Baues-Wirsching cohomology, Thomason cohomology, Grothendieck construction, Quillen’s Theorem A
2020 Mathematics Subject Classification. Primary: 18G90; Secondary: 55U10, 18G30, 18G35, 18G40
Corresponding Author: Ergün Yalçın, Email: yalcine@fen.bilkent.edu.tr
Address: Bilkent University, Department of Mathematics, 06800, Bilkent, Ankara, Türkiye
Contributing Author: Mehmet Kırtışoğlu, Email: m.kirtisoglu@bilkent.edu.tr
Address: Bilkent University, Department of Mathematics, 06800, Bilkent, Ankara, Türkiye
1. Introduction and statement of results
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor between two small categories. For each $d\in\mathcal{D}$, let $\varphi/d$ denote the comma category with objects given by the pairs $(c,\mu)$ where $c\in\mathcal{C}$ and $\mu:\varphi(c)\to d$ is a morphism in $\mathcal{D}$. The assignment $d\to\varphi/d$ for every $d\in\mathcal{D}$ defines a functor $\varphi/-:\mathcal{D}\to Cat$. Composing this functor with the nerve functor we obtain a functor $N(\varphi/-):\mathcal{D}\to sSet$. We define the homotopy colimit $\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)$ as the diagonal of the simplicial replacement of the functor $N(\varphi/-)$.
Definition 1.1.
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor. For each $d\in\mathcal{D}$, there is a simplicial map $j_{d}:N(\varphi/d)\to N\mathcal{C}$ induced by the functor $\varphi/d\to\mathcal{C}$ that sends the pair $(c,\mu)$ in $\varphi/d$ to $c\in\mathcal{C}$. The simplicial maps $j_{d}$ induce a simplicial map
$\kappa:\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\to N\mathcal{C}$
defined by $\kappa(\sigma,\tau,\mu)=\tau$ (see Lemma 5.1).
More details on the above definitions can be found in Section 5. It is shown in [16, Theorem A] that $\kappa$ is a homotopy equivalence. This homotopy equivalence is the first step in the proof of Quillen’s Theorem A. In this paper we prove that the homotopy equivalence $\kappa$ induces an isomorphism for the cohomology of the associated simplicial sets with coefficients in most general coefficient systems.
For a small category $\mathcal{C}$, there are different notions of the cohomology of $\mathcal{C}$ depending on the type of the coefficient system that is chosen. Let $R$ be a commutative ring, then a functor $M:\mathcal{C}\to R$-Mod is called an $R\mathcal{C}$-module. The constant functor with value $R$ is denoted by $\underline{R}$. The cohomology $H^{*}(\mathcal{C};M)$ of a small category $\mathcal{C}$ with coefficients in an $R\mathcal{C}$-module $M$ is defined as the ext-group $\mathrm{Ext}_{R\mathcal{C}}(\underline{R};M)$ over the module category of $R\mathcal{C}$-modules (see Section 2.1). There is a standard resolution for $\underline{R}$ as an $R\mathcal{C}$-module that allows us to express the cohomology groups $H^{*}(\mathcal{C};M)$ as cohomology groups of an explicit cochain complex. We refer to this cohomology theory as the Quillen cohomology of a small category.
Given a small category $\mathcal{C}$, the category of factorizations $\mathfrak{F}\mathcal{C}$ is the category whose objects are the morphisms in $\mathcal{C}$ and whose morphisms $f\to g$ are given by a pair of morphisms $(u,v)$ in $\mathcal{C}$ such that $ufv=g$. A functor $M:\mathfrak{F}\mathcal{C}\to R$-Mod is called a natural system for $\mathcal{C}$ over $R$. The Baues-Wirsching cohomology $H^{*}_{BW}(\mathcal{C};M)$ of the category $\mathcal{C}$ with coefficients in a natural system $M$ is either defined using an explicit cochain complex or defined as the ext-group $\mathrm{Ext}_{R\mathfrak{F}\mathcal{C}}(\underline{R},M)$ (see Section 2.2).
For a small category $\mathcal{C}$, let $N\mathcal{C}$ denote the nerve of $\mathcal{C}$. A functor $\mathcal{M}:\Delta(N\mathcal{C})\to R$-Mod is called a coefficient system for $N\mathcal{C}$ where $\Delta(N\mathcal{C})$ denotes the category of simplices of $N\mathcal{C}$ (see Definition 3.1). For a coefficient system $\mathcal{M}$ for $N\mathcal{C}$, the Thomason cohomology $H^{*}_{Th}(\mathcal{C};\mathcal{M})$ is defined as the Quillen cohomology of the category $\Delta(N\mathcal{C})$ with coefficients in $\mathcal{M}$ (see Section 3.4). The cohomology groups $H^{*}(\mathcal{C};M)$ and $H^{*}_{BW}(\mathcal{C};M)$ can be realized as the Thomason cohomology $H^{*}_{Th}(\mathcal{C};\mathcal{M})$ of $\mathcal{C}$ with suitable choices of coefficient systems $\mathcal{M}$ (see Lemma 3.7).
The main theorem of the paper is the following:
Theorem 1.2.
Let $\varphi:\mathcal{C}\rightarrow\mathcal{D}$ be a functor between two small categories, and let $X=\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)$ and $\kappa:X\to N\mathcal{C}$ be the homotopy equivalence defined in Definition 1.1. Then for every coefficient system $\mathcal{M}:\Delta(N\mathcal{C})\rightarrow R$-Mod, the homomorphism
$\kappa^{*}:H^{*}_{Th}(\mathcal{C};\mathcal{M})\smash{\,\mathop{\longrightarrow}\limits^{\cong}\,}H^{*}(X;\kappa^{*}\mathcal{M})$
induced by $\kappa$ is an isomorphism.
Theorem 1.2 is proved by constructing a double complex of projective $R\Delta(N\mathcal{C})$-modules whose total complex gives a projective resolution for the constant functor $\underline{R}$ as an $R\Delta(N\mathcal{C})$-module. This idea of constructing a projective resolution as a double complex was also used by Cegarra in the proof of [3, Theorem 1]. As a special case of Theorem 1.2, we obtain that the simplicial map $\kappa$ gives isomorphisms for the Baues-Wirsching and Quillen cohomologies of $\mathcal{C}$.
As a consequence of Theorem 1.2, we also prove a version of Quillen’s Theorem A for Thomason cohomology. Let
$U:\varphi/-\to\mathrm{id}/$-
be the natural transformation defined by sending $d\in\mathcal{D}$ to the functor $U(d):\varphi/d\to\mathrm{id}/d:(c,\mu)\to(\varphi(c),\mu)$. For every $d\in\mathcal{D}$, there is a commuting diagram of simplicial maps
${N(\varphi/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}}$$\scriptstyle{NU(d)}$${N\mathcal{C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{N\varphi}$${N(\mathrm{id}/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}^{\prime}}$${N\mathcal{D}}$
where $j_{d}$ and $j_{d}^{\prime}$ are the simplicial maps defined in Definition 1.1.
Theorem 1.3.
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor and $\mathcal{M}$ be any coefficient system for $N\mathcal{D}$. Suppose that for each $d\in\mathcal{D}$, the simplicial map $NU(d):N(\varphi/d)\to N(\mathrm{id}/d)$ induces an isomorphism
$H^{*}(N(\mathrm{id}/d);\mathcal{M}_{d})\cong H^{*}(N(\varphi/d);NU(d)^{*}\mathcal{M}_{d})$
where $\mathcal{M}_{d}=(j_{d}^{\prime})^{*}\mathcal{M}$. Then the map
$\varphi^{*}:H^{*}_{Th}(\mathcal{D};\mathcal{M})\to H^{*}_{Th}(\mathcal{C};\varphi^{*}\mathcal{M})$
induced by $\varphi$ is an isomorphism.
To prove Theorem 1.3, we construct a spectral sequence that converges to the Thomason cohomology of the category $\mathcal{C}$. As a special case of this spectral sequence we obtain a spectral sequence for the Thomason cohomology of the Grothendieck construction $\int_{\mathcal{D}}F$ of a functor $F:\mathcal{D}\to Cat$.
Theorem 1.4.
Let $F:\mathcal{D}\to Cat$ be a functor and $\int_{\mathcal{D}}F$ be the Grothendieck construction of $F$. Let $\pi:\int_{\mathcal{D}}F\to\mathcal{D}$ be the canonical functor that sends a pair $(d,x)\in\int_{\mathcal{D}}F$ to $d\in\mathcal{D}$, and $\widetilde{F}:\mathcal{D}\to Cat$ be the functor that sends $d\in\mathcal{D}$ to the comma category $\pi/d$. Then for every coefficient system $\mathcal{M}$ defined on $N(\int_{\mathcal{D}}F)$, there is a spectral sequence
$E_{2}^{p,q}=H^{p}_{Th}(\mathcal{D};\mathcal{H}^{q}_{\mathcal{M}})\Rightarrow H^{p+q}_{Th}(\textstyle\int_{\mathcal{D}}F;\mathcal{M})$
where $\mathcal{H}^{q}_{\mathcal{M}}$ is the coefficient system on $N\mathcal{D}$ that sends a simplex $\sigma=(d_{0}\to\cdots\to d_{p})\in N\mathcal{D}_{p}$ to the cohomology module $H^{q}(N\widetilde{F}(d_{0});j_{d_{0}}^{*}\mathcal{M})$.
A similar spectral sequence that converges to the Baues-Wirsching cohomology of the Grothendieck construction $\int_{\mathcal{D}}F$ was constructed earlier by Pirashvili and Redondo [15] and by Cegarra [3] using different methods under certain conditions on the coefficient system $\mathcal{M}$. Theorem 1.4 can be considered as a generalization of these spectral sequences to the Thomason cohomology of small categories.
Notation and Conventions: All the categories are assumed to be small (or equivalent to a small category). We write $d\in\mathcal{D}$ when $d$ is an object of the category $\mathcal{D}$.
2. Modules and cohomology for small categories
Let $\mathcal{C}$ be a small category and $R$ a commutative ring. A functor $M:\mathcal{C}\to R$-Mod is called a (covariant) $R\mathcal{C}$-module. The category of $R\mathcal{C}$-modules is an abelian category, which has enough projective and injective modules. The exactness of a short exact sequence $M_{1}\to M_{2}\to M_{3}$ in $R\mathcal{C}$-Mod is defined by component-wise exactness of the corresponding $R$-modules $M_{1}(x)\to M_{2}(x)\to M_{3}(x)$.
There is an explicit construction of projective $R\mathcal{C}$-modules using the Yoneda Lemma: For every $x\in\mathcal{C}$, let $R\mathrm{Mor}_{\mathcal{C}}(x,?)$ denote the functor $\mathcal{C}\to R$-Mod that sends an object $c\in\mathcal{C}$ to the free $R$-module with basis $\mathrm{Mor}_{\mathcal{C}}(x,c)$. By the Yoneda Lemma, for every $R\mathcal{C}$-module $M$
$\mathrm{Hom}_{R\mathcal{C}}(R\mathrm{Mor}_{\mathcal{C}}(x,?),M)\cong M(x).$
Using this isomorphism, one can show that the $R\mathcal{C}$-module $R\mathrm{Mor}_{\mathcal{C}}(x,?)$ is projective (see [14], [20] for details). Having enough projective modules allows us to construct projective resolutions and define ext-groups over $R\mathcal{C}$-modules.
Definition 2.1.
For every pair of $R\mathcal{C}$-modules $M$ and $N$, the ext-group $\mathrm{Ext}^{n}_{R\mathcal{C}}(N,M)$ is the $n$-th cohomology of the cochain complex $\mathrm{Hom}_{R\mathcal{C}}(P_{*},M)$ where $P_{*}\to N$ is any projective resolution of $N$ as an $R\mathcal{C}$-module.
Instead of covariant functors $M:\mathcal{C}\to R$-Mod, one can consider the abelian category of contravariant functors $M:\mathcal{C}^{op}\to R$-Mod. We call such a module a contravariant $R\mathcal{C}$-module (or a right $R\mathcal{C}$-module). The contravariant $R\mathcal{C}$-modules defined by $R\mathrm{Mor}_{\mathcal{C}}(?,x)$ for $x\in\mathcal{C}$ are projective, and the ext-groups $\mathrm{Ext}^{*}_{R\mathcal{C}}(M,N)$ for contravariant $R\mathcal{C}$-modules $M$ and $N$ are defined in a similar way using the resolutions of contravariant $R\mathcal{C}$-modules.
2.1. Cohomology of small categories
For a small category $\mathcal{C}$ and a commutative ring $R$, the constant functor $\underline{R}$ is the $R\mathcal{C}$-module such that for each $c\in\mathcal{C}$, $\underline{R}(c)=R$, and for every morphism $\varphi:c\to c^{\prime}$ in $\mathcal{C}$, the induced map $\underline{R}(\varphi):R\to R$ is the identity map.
Definition 2.2.
The (Quillen) cohomology of a small category $\mathcal{C}$ with coefficients in an $R\mathcal{C}$-module $M$ is defined by
$H^{*}(\mathcal{C};M):=\mathrm{Ext}^{*}_{R\mathcal{C}}(\underline{R},M).$
Remark 2.3.
There is no standard terminology for the above definition of the cohomology of a small category. Throughout the paper we refer this cohomology definition the Quillen cohomology of a small category to distinguish it from the Baues-Wirsching and the Thomason cohomology of a small category.
The nerve $N\mathcal{C}$ of the category $\mathcal{C}$ is a simplicial set whose $n$-simplices, $N\mathcal{C}_{n}$, are given by the length $n$ chains of composable morphisms $c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n}$ in $\mathcal{C}$. For $n=0$, we take $NC_{0}$ to be the set of objects in $\mathcal{C}$. For $n\geq 1$, the boundary map $d_{i}:N\mathcal{C}_{n}\to N\mathcal{C}_{n-1}$, $0\leq i\leq n$, is defined by
$d_{i}(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})=\begin{cases}c_{1}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{2}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n}&\text{ if }i=0\\ c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots c_{i-1}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{i+1}\alpha_{i}}\,}c_{i+1}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n}&\text{ if }0<i<n\\ c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n-1}}\,}c_{n-1}&\text{ if }i=n\end{cases}$
The degeneracy map $s_{i}:N\mathcal{C}_{n}\to N\mathcal{C}_{n+1}$, $0\leq i\leq n$, is defined by
$s_{i}(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})=c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots c_{i}\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}c_{i}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n}.$
For each $n\geq 0$, let
$P_{n}=\bigoplus_{\sigma=(c_{0}\to\cdots\to c_{n})\in N\mathcal{C}_{n}}R\mathrm{Mor}_{\mathcal{C}}(c_{n},?),$
and for every $c\in\mathcal{C}$, let $\partial_{n}(c):P_{n}(c)\to P_{n-1}(c)$ be the map defined by
$\partial_{n}(\sigma,\alpha)=\sum_{i=0}^{n-1}(-1)^{i}(d_{i}\sigma,\alpha)+(-1)^{n}(d_{n}\sigma,\alpha\circ\alpha_{n})$
where $\sigma=(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$ is a simplex in $N\mathcal{C}_{n}$ and $\alpha:c_{n}\to c$ is a morphism in $\mathcal{C}$.
Lemma 2.4.
The chain complex
$\widetilde{P}_{*}:\cdots\to P_{n}\smash{\,\mathop{\longrightarrow}\limits^{\partial_{n}}\,}P_{n-1}\to\cdots\to P_{0}\smash{\,\mathop{\longrightarrow}\limits^{\epsilon}\,}\underline{R}\to 0$
is a projective resolution of $\underline{R}$ as an $R\mathcal{C}$-module. This resolution is called the standard resolution for $\mathcal{C}$.
Proof.
For each $c\in\mathcal{C}$, the chain complex $\widetilde{P}_{*}(c)$ is exact because there is a splitting $s_{*}:P_{*}(c)\to P_{*+1}(c)$ defined by
$s_{n}((c_{0}\to\cdots\to c_{n}),\alpha:c_{n}\to c)=(-1)^{n+1}((c_{0}\to\cdots\to c_{n}\smash{\,\mathop{\longrightarrow}\limits^{\alpha}\,}c),\mathrm{id}:c\to c)$
for $n\geq 0$, and $s_{-1}(1)=(c,\mathrm{id}:c\to c)$. Hence $P_{*}$ is a projective resolution of $\underline{R}$. ∎
Using the standard resolution, one can describe the cohomology of a category $\mathcal{C}$ with coefficients in an $R\mathcal{C}$-module $M$ as the cohomology of the cochain complex $C^{*}(\mathcal{C};M)$ where for each $n\geq 0$,
$C^{n}(\mathcal{C};M)=\{f:N\mathcal{C}_{n}\to\coprod_{c\in\mathcal{C}}M(c)\,|\ f(c_{0}\to\cdots\to c_{n})\in M(c_{n})\}$
and the coboundary map $\delta^{n-1}:C^{n-1}(\mathcal{C};M)\to C^{n}(\mathcal{C};M)$ is defined by
$(\delta^{n-1}f)(\sigma)=\sum_{i=0}^{n-1}(-1)^{i}f(d_{i}\sigma)+(-1)^{n}M(\alpha_{n})f(d_{n}\sigma)$
for every $\sigma=(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$ in $N\mathcal{C}_{n}$. For more details on cohomology of categories we refer the reader to [16], [17], and [21].
2.2. Baues-Wirsching Cohomology
The Baues-Wirsching cohomology [1] of a small category $\mathcal{C}$ is a cohomology theory for $\mathcal{C}$ with more general coefficients, called natural systems. Throughout the section, $R$ denotes a commutative ring with unity.
Definition 2.5.
The category of factorizations of a small category $\mathcal{C}$ is the category $\mathfrak{F}\mathcal{C}$ whose objects are the morphisms $\alpha:x\to y$ in $\mathcal{C}$, and whose morphisms $(u,v):\alpha\to\alpha^{\prime}$ are given by a pair of morphisms in $\mathcal{C}$ such that the following diagram commutes:
[…]
i.e., $\alpha^{\prime}=u\alpha v$. The composition $(u^{\prime},v^{\prime})\circ(u,v)$ of two morphisms $(u,v):\alpha\to\alpha^{\prime}$ and $(u^{\prime},v^{\prime}):\alpha^{\prime}\to\alpha^{\prime\prime}$ is defined by $(u^{\prime}u,vv^{\prime}):\alpha\to\alpha^{\prime\prime}$. A functor $M:\mathfrak{F}\mathcal{C}\to R$-Mod is called a natural system for $\mathcal{C}$ over $R$.
For the morphism $(u,1):\alpha\to u\alpha$ in $\mathfrak{F}\mathcal{C}$, the induced map $M(u,1):M(\alpha)\to M(u\alpha)$ is denoted $u_{*}$, and for $(1,v):\alpha\to\alpha v$ in $\mathfrak{F}\mathcal{C}$, the induced map $M(1,v):M(\alpha)\to M(\alpha v)$ is denoted $v^{*}$. So, for each morphism $(u,v)$ in $\mathfrak{F}\mathcal{C}$, we have $M(u,v)=u_{*}v^{*}=v^{*}u_{*}$.
Definition 2.6.
Let $\mathcal{C}$ be a small category and $M$ be a natural system for $\mathcal{C}$ over $R$ . The Baues-Wirsching cohomology $H^{*}_{BW}(\mathcal{C};M)$ of $\mathcal{C}$ with coefficients in $M$ is the cohomology of a cochain complex $C^{*}_{BW}(\mathcal{C};M)$ where for all $n\geq 0$, $C^{n}_{BW}(\mathcal{C};M)$ is the abelian group of all functions
$f:N\mathcal{C}_{n}\to\coprod_{\alpha\in\mathrm{Mor}(\mathcal{C})}M(\alpha)$
such that $f(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})\in M(\alpha_{n}\cdots\alpha_{1})$ if $n>0$. For $n=0$, we require $f(c)\in M(\mathrm{id}_{c})$ for all $c\in\mathcal{C}$. For $n>1$, the coboundary map
$\delta^{n-1}:C^{n-1}_{BW}(\mathcal{C};M)\to C^{n}_{BW}(\mathcal{C};M)$
is defined by
$\delta^{n-1}(f)(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$ $=\alpha_{1}^{*}f(c_{1}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{2}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$
$+\sum_{i=1}^{n-1}(-1)^{i}f(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots c_{i-1}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{i+1}\alpha_{i}}\,}c_{i+1}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$
$+(-1)^{n}(\alpha_{n})_{*}f(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n-1}}\,}c_{n-1}).$
For $n=1$, we have $\delta^{0}(f)(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha}\,}c_{1})=\alpha^{*}f(c_{1})-\alpha_{*}f(c_{0})$.
The Baues-Wirsching cohomology satisfies many properties that one would expect to hold for a cohomology theory. For example, if $\varphi:\mathcal{C}\to\mathcal{D}$ is an equivalence of categories then the induced map
$\varphi^{*}:H^{*}_{BW}(\mathcal{D};M)\to H^{*}_{BW}(\mathcal{C};\varphi^{*}M)$
is an isomorphism for every natural system $M:\mathfrak{F}\mathcal{D}\to R$-Mod (see [1, Theorem 1.11]).
Similar to the Quillen cohomology of small categories $H^{*}(\mathcal{C};M)$, Baues-Wirsching cohomology can also be interpreted as Ext-groups over a category of modules.
Proposition 2.7.
[1, Theorem 4.4] Let $\underline{R}:\mathfrak{F}\mathcal{C}\to R$-Mod denote the constant functor for $F\mathcal{C}$ over $R$. For every natural system $M$ for $\mathcal{C}$, and for $n\geq 0$, there is an isomorphism
$H^{n}_{BW}(\mathcal{C};M)\cong\mathrm{Ext}^{n}_{R\mathfrak{F}\mathcal{C}}(\underline{R},M)$
which is natural in $M$.
To prove Proposition 2.7, Baues and Wirsching construct a projective resolution $P_{*}\to\underline{R}$ of the constant functor $\underline{R}$ as a $R\mathfrak{F}\mathcal{C}$-module such that $\mathrm{Hom}_{R\mathfrak{F}\mathcal{C}}(P_{*},M)$ is a cochain complex isomorphic to $C^{*}_{BW}(\mathcal{C};M)$. Note that for every object $\alpha\in\mathfrak{F}\mathcal{C}$, the $R\mathfrak{F}\mathcal{C}$-module $R\mathrm{Mor}_{\mathfrak{F}\mathcal{C}}(\alpha,?)$ is projective. One can use these modules to construct the projective resolution (see [1, Theorem 4.4] for details).
The Baues-Wirsching cohomology generalizes the Quillen cohomology of a small category in the following sense: For a small category $\mathcal{C}$, there are functors $S:\mathfrak{F}\mathcal{C}\to\mathcal{C}^{op}$ and $T:\mathfrak{F}\mathcal{C}\to\mathcal{C}$ defined by $S(\alpha:x\to y)=x$ and $T(\alpha:x\to y)=y$ on objects. A morphism $(u,v):\alpha\to\alpha^{\prime}$ in $\mathfrak{F}\mathcal{C}$ is taken to $v:x^{\prime}\to x$ and $u:y\to y^{\prime}$ under the functors $S$ and $T$, respectively.
Lemma 2.8 (Baues-Wirsching [1]).
(i) If $M$ is a natural system defined by the composition
$M:\mathfrak{F}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{T}\,}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{N}\,}R\text{-mod},$
then $H^{*}_{BW}(\mathcal{C};M)$ is isomorphic to the cohomology $H^{*}(\mathcal{C};N)$ of $\mathcal{C}$ with coefficients in the (covariant) $R\mathcal{C}$-module $N$.
(ii) If $M$ is a natural system defined by the composition
$M:\mathfrak{F}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{S}\,}\mathcal{C}^{op}\smash{\,\mathop{\longrightarrow}\limits^{N^{\prime}}\,}R\text{-mod},$
then $H^{*}_{BW}(\mathcal{C};M)$ is isomorphic to the cohomology $H^{*}(\mathcal{C};N^{\prime})$ of $\mathcal{C}$ with coefficients in the contravariant $R\mathcal{C}$-module $N^{\prime}$.
3. Cohomology of simplicial sets and Thomason Cohomology
In this section we give definitions for cohomology of simplicial sets with general coefficient systems and define Thomason cohomology for small categories. For more details on this material we refer the reader to [5], [7] and [8].
3.1. Coefficient systems for simplicial sets
The simplex category $\Delta$ is the category whose objects are ordered finite sets $n]=\{0,1,\dots,n\}$, and whose morphisms are given by weakly order-preserving functions between them. Morphisms in the simplex category are generated by coboundary maps $d^{i}:[n-1]\to[n]$, $0\leq i\leq n$, and codegeneracy maps $s^{i}:[n+1]\to[n]$, $0\leq i\leq n$, which satisfy certain relations (see [9]).
A simplicial object in the category $\mathcal{C}$ is a functor $\Delta^{op}\to\mathcal{C}$. A simplicial object in sets is called a simplicial set. We write $X_{n}$ for $X([n])$. For a morphism $f:[m]\to[n]$, we denote the induced map $X(f):X[n]\to X[m$] by $f^{*}:X_{n}\to X_{m}$. The induced maps $(d^{i})^{*}:X_{n}\to X_{n-1}$ and $(s^{i})^{*}:X_{n}\to X_{n+1}$ are denoted $d_{i}$ and $s_{i}$, respectively, and called the boundary and degeneracy maps.
A natural transformation $X\to Y$ between two simplicial sets $X,Y$ is called a simplicial map. The category of simplicial sets is denoted $sSet$.
Definition 3.1.
Given a simplicial set $X$, the category of simplices of $X$, or the category of elements of $X$, is the category $\Delta(X)$ whose objects are the pairs $([n],\sigma)$ where $\sigma\in X_{n}$, and the morphisms $f:([m],\tau)\to([n],\sigma)$ in $\Delta(X)$ are given by the morphisms $f:[m]\to[n]$ in $\Delta$ such that $f^{*}(\sigma)=\tau$. A functor $\mathcal{M}:\Delta(X)\to R$-Mod is called a coefficient system $\mathcal{M}$ for the simplicial set $X$ over the commutative ring $R$.
We write $\mathcal{M}(\sigma)$ for $\mathcal{M}([n],\sigma)$, and for each morphism $f:([m],\tau)\to([n],\sigma)$, the induced map $\mathcal{M}(f):\mathcal{M}(\tau)\to\mathcal{M}(\sigma)$ is denoted $f_{*}$. In particular, for each $i$, the coboundary and codegeneracy maps induce $R$-module homomorphisms
$d^{i}_{*}:\mathcal{M}(d_{i}\sigma)\to\mathcal{M}(\sigma)\quad\text{ and }\quad s^{i}_{*}:\mathcal{M}(s_{i}\sigma)\to\mathcal{M}(\sigma).$
A coefficient system is called a local coefficient system if for every morphism $f:([m],\tau)\to([n],\sigma)$ in $\Delta(X)$, the induced homomorphism $f_{*}:\mathcal{M}(\tau)\to\mathcal{M}(\sigma)$ is an isomorphism.
3.2. Cohomology of simplicial sets with general coefficients
Definition 3.2.
Let $X$ be a simplicial set and $\mathcal{M}$ a coefficient system for $X$. Consider the cochain complex $C^{*}(X;\mathcal{M})$ such that for each $n\geq 0$,
$C^{n}(X;\mathcal{M})=\{f:X_{n}\to\prod_{\sigma\in X_{n}}\mathcal{M}(\sigma)\,|\,f(\sigma)\in\mathcal{M}(\sigma)\}$
where the coboundary map $\delta^{n-1}:C^{n-1}(X;\mathcal{M})\to C^{n}(X;\mathcal{M})$ is given by
$\delta^{n-1}(f)(\sigma)=\sum_{i=0}^{n}(-1)^{i}d^{i}_{*}\bigl(f(d_{i}\sigma)\bigr)$
for every $f\in C^{n-1}(X;\mathcal{M})$ and $\sigma\in X_{n}$. The cohomology $H^{*}(X;\mathcal{M})$ of the simplicial set $X$ with coefficients in $\mathcal{M}$ is the cohomology of the cochain complex $C^{*}(X;M)$.
An alternative way to define cohomology of a simplicial set $X$ with coefficients in $\mathcal{M}$ is to define it as the cohomology of the category $\Delta(X)$ with coefficients in $\mathcal{M}$ by considering $\mathcal{M}$ as a covariant $R\Delta(X)$-module. It is a well-known fact that these two cohomology definitions are isomorphic (see Gabriel–Zisman [5, Appendix II]).
Proposition 3.3.
Let $X$ be a simplicial set and $\mathcal{M}$ be a coefficient system for $X$ over $R$. The cohomology groups $H^{*}(X;\mathcal{M})$, defined as the cohomology of the cochain complex $C^{*}(X;\mathcal{M})$, are isomorphic to the cohomology groups $H^{*}(\Delta(X);\mathcal{M})$ of the category $\Delta(X)$ with coefficients in the covariant module $\mathcal{M}:\Delta(X)\to R$-Mod.
Proof.
The proof is similar to earlier proofs of this type. We show that there is a projective resolution $P_{*}$ of $\underline{R}$ as an $R\Delta(X)$-module such that $\mathrm{Hom}_{R\Delta(X)}(P_{*},\mathcal{M})$ is isomorphic to the cochain complex $C^{*}(X;\mathcal{M})$. For each $q\geq 0$, let
$P_{q}=\bigoplus_{\sigma\in X_{q}}R\mathrm{Mor}_{\Delta(X)}(\sigma,?).$
For each simplex $\theta\in X_{n}$, $P_{q}(\theta)$ is a free $R$-module with basis given by
$\mathcal{B}_{q}(\theta)=\{(\sigma,f)\,|\,\sigma\in X_{q},f:[q]\to[n]\text{ such that }f^{*}(\theta)=\sigma\}.$
The boundary map $\partial_{q}:P_{q}(\theta)\to P_{q-1}(\theta)$ is defined by
$\partial_{q}(\sigma,f)=\sum_{i=0}^{q}(-1)^{i}(d_{i}\sigma,f\circ d^{i}).$
For each $n\geq 0$, let $\Delta[n]$ denote the simplicial set whose $k$-simplices are given by $(\Delta[n])_{k}=\mathrm{Mor}_{\Delta}([k],[n])$. Note that for every $\theta\in X_{n}$, the set $\mathcal{B}_{q}(\theta)$ is isomorphic to $\mathrm{Mor}_{\Delta}([q],[n])$. Hence the homology of the chain complex
$\widetilde{P}_{*}:\cdots\to P_{q}(\theta)\smash{\,\mathop{\longrightarrow}\limits^{\partial_{q}}\,}P_{q-1}(\theta)\to\cdots\to P_{0}(\theta)\to R\to 0$
is isomorphic to the reduced homology of simplicial set $\Delta[n]$. Since $|\Delta[n]|=\Delta^{n}$ is contractible, this reduced homology is trivial. Hence $P_{*}\to\underline{R}$ is a projective resolution of $\underline{R}$ as an $R\Delta(X)$-module. It is easy to see that for every $R\Delta(X)$-module $\mathcal{M}$, the cochain complex $\mathrm{Hom}_{R\Delta(X)}(P_{*},\mathcal{M})$ is isomorphic to $C^{*}(X;\mathcal{M})$. ∎
3.3. Induced Maps on Cohomology
Let $\lambda:X\to Y$ be a simplicial map between two simplicial sets. The map $\lambda$ induces a functor $\Delta(\lambda):\Delta(X)\to\Delta(Y)$. For every coefficient system $\mathcal{M}:\Delta(Y)\to R$-Mod for $Y$, the composition
$\lambda^{*}\mathcal{M}:\Delta(X)\xrightarrow{\Delta(\lambda)}\Delta(Y)\xrightarrow{\mathcal{M}}R\text{-Mod}$
is a coefficient system for $X$.
Lemma 3.4.
The simplicial map $\lambda:X\to Y$ induces an $R$-module homomorphism
$\lambda^{*}:H^{*}(Y;\mathcal{M})\to H^{*}(X;\lambda^{*}\mathcal{M})$
for every coefficient system $\mathcal{M}$ for $Y$.
Proof.
The simplicial map $\lambda$ induces a chain map
$\lambda^{*}:C^{n}(Y,\mathcal{M})\to C^{n}(X;\lambda^{*}\mathcal{M})$
defined by $\lambda^{*}(f)(\sigma)=f(\lambda(\sigma))$ for every simplex $\sigma$ in $X$. Note that
$\lambda^{*}(f)(\sigma)=f(\lambda(\sigma))\in\mathcal{M}(\lambda(\sigma))=(\lambda^{*}\mathcal{M})(\sigma).$
The chain map $\lambda^{*}$ induces the desired homomorphism between the corresponding cohomology modules. ∎
Definition 3.5.
The category of simplicial sets with coefficients is defined as the category whose objects are the pairs $(X,\mathcal{M})$ where $X$ is a simplicial set and $\mathcal{M}$ is a coefficient system on $X$. A morphism of spaces of coefficients $(X,\mathcal{M}_{X})\to(Y,\mathcal{M}_{Y})$ is a pair of maps $(\lambda,\theta)$ where $\lambda:X\to Y$ is a simplicial map and $\theta:\lambda^{*}\mathcal{M}_{Y}\to\mathcal{M}_{X}$ is a morphism of $\Delta(X)$-modules.
A morphism of pairs $(\lambda,\theta):(X,\mathcal{M}_{X})\to(Y,\mathcal{M}_{Y})$ induces a homomorphism of cohomology groups
$(\lambda,\theta)^{*}:H^{*}(Y,\mathcal{M}_{Y})\smash{\,\mathop{\longrightarrow}\limits^{\lambda^{*}}\,}H^{*}(X;\lambda^{*}\mathcal{M}_{Y})\smash{\,\mathop{\longrightarrow}\limits^{\theta_{*}}\,}H^{*}(X;\mathcal{M}_{X}).$
3.4. The Thomason cohomology of a category
Definition 3.6.
Let $\mathcal{C}$ be a small category. A coefficient system for the simplicial set $N\mathcal{C}$ is called a Thomason natural system for $\mathcal{C}$. Given a Thomason system $\mathcal{M}:\Delta(N\mathcal{C})\to R$-mod, the cohomology of the simplicial set $N\mathcal{C}$ with coefficients in $\mathcal{M}$ is called the Thomason cohomology of $\mathcal{C}$ and is denoted
$H^{*}_{Th}(\mathcal{C};\mathcal{M}):=H^{*}(N\mathcal{C};\mathcal{M}).$
The Thomason cohomology of a category was first introduced by Galvez-Carrillo, Neumann, and Tonks in [7]. Thomason cohomology generalizes the Baues-Wirsching cohomology of a category in the following sense:
Let $\chi:\Delta(N\mathcal{C})\to\mathfrak{F}\mathcal{C}$ be the functor that sends $\sigma=(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$ in $N\mathcal{C}_{n}$ to $\sigma(0,n):=\alpha_{n}\cdots\alpha_{1}:c_{0}\to c_{n}$ in $\mathfrak{F}\mathcal{C}$. For each morphism $f:([m],\tau)\to([n],\sigma)$ in $\Delta(N\mathcal{C})$, where $\tau=(c_{0}^{\prime}\smash{\,\mathop{\longrightarrow}\limits^{\beta_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\beta_{m}}\,}c_{m}^{\prime})$ and $\sigma=(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$, we define $\chi(f)$ to be the morphism
$(\alpha_{n}\cdots\alpha_{f(m)+1},\alpha_{f(0)}\cdots\alpha_{1}):\beta_{m}\cdots\beta_{1}\to\alpha_{n}\cdots\alpha_{1}.$
Lemma 3.7 ([7, Thm 2.1]).
If $\mathcal{M}$ is a coefficient system for $N\mathcal{C}$ defined by the composition
$\mathcal{M}:\Delta(N\mathcal{C})\smash{\,\mathop{\longrightarrow}\limits^{\chi}\,}\mathfrak{F}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{M}\,}R\text{-Mod}$
for some natural system $M$, then we have
$H^{*}_{Th}(\mathcal{C};\mathcal{M})\cong H^{*}_{BW}(\mathcal{C};M).$
4. Bisimplicial objects and the Dold-Puppe theorem
4.1. Homotopy colimits
Let $\Delta$ denote the simplex category. A bisimplicial object in the category $\mathcal{C}$ is a functor $X:\Delta^{op}\times\Delta^{op}\to\mathcal{C}$. We denote the object $X([p]\times[q])$ by $X_{p,q}$. The simplicial objects $X_{p,*}$ and $X_{*,q}$ are called vertical and horizontal simplicial objects associated to $X$. The diagonal of a bisimplicial object is the simplicial object defined by
$\mathrm{diag}X:\Delta^{op}\to\Delta^{op}\times\Delta^{op}\to\mathcal{C}$
where the first functor is the diagonal functor defined by $n]\to([n],[n])$.
A bisimplicial set is a bisimplicial object in sets. We define homotopy colimits of simplicial sets using the notation and terminology in [13]. Let $\mathcal{D}$ be a small category and $F:\mathcal{D}\to sSet$ a functor. Consider the bisimplicial set $N(\mathcal{D};F)$ such that for $p,q\geq 0$,
$N(\mathcal{D};F)_{p,q}=\coprod_{\sigma=(d_{0}\to\dots\to d_{p})\in N\mathcal{D}_{p}}F(d_{0})_{q}$
with vertical and horizontal induced maps are defined as follows: If $f_{h}\times f_{v}:([p^{\prime}],[q^{\prime}])\to([p],[q])$ is a morphism in $\Delta\times\Delta$, then for every $(\sigma,\tau)$ with
$\sigma=(d_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}d_{1}\to\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{p}}\,}d_{p})\in N\mathcal{D}_{p}\quad\text{ and }\quad\tau=(x_{0}\smash{\,\mathop{\longrightarrow}\limits^{\omega_{1}}\,}x_{1}\to\cdots\smash{\,\mathop{\longrightarrow}\limits^{\omega_{q}}\,}x_{q})\in F(d_{0})_{q},$
we define
$(f_{h},f_{v})^{*}(\sigma,\tau)=(f_{h}^{*}(\sigma),F(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))).$ (1)
This definition respects the composition of functions in $\Delta\times\Delta$.
Definition 4.1.
Let $F:\mathcal{D}\to sSet$ be a functor. The bisimplicial set $N(\mathcal{D};F)$ is called the simplicial replacement of $F$. We define the homotopy colimit $\operatorname*{hocolim}_{\mathcal{D}}F$ to be the diagonal $\mathrm{diag}N(\mathcal{D};F)$ of the simplicial replacement of $F$ (see [2, §12.5.2]).
4.2. The Dold-Puppe Theorem
A simplicial object in $R$-modules is called a simplicial $R$-module. A functor $A:\Delta^{op}\times\Delta^{op}\to R$-Mod is called a bisimplicial $R$-module.
Definition 4.2.
The Moore complex of a simplicial $R$-module AA is the chain complex whose $n$-chains are given by $A_{n}$ and whose boundary maps $\partial_{n}:A_{n}\to A_{n-1}$ are defined by $\partial_{n}(x)=\sum_{i=0}^{n}(-1)^{i}d_{i}(x)$. The Moore complex of $A$, denoted by $A_{*}$, is a chain complex and its homology is denoted by $H_{*}(A)$.
For a bisimplicial $R$-module $A:\Delta^{op}\times\Delta^{op}\to R$-Mod, we define the associated double complex $A_{*,*}$ to be the double complex with horizontal and vertical boundary maps
$\partial^{h}_{p}=\sum_{i=0}^{p}(-1)^{i}d_{i}^{h}\ \text{ and }\ \partial^{v}_{q}=\sum_{j=0}^{q}(-1)^{j}d_{j}^{v}.$
The total complex of the double complex $A_{*,*}$ is the chain complex $\mathrm{Tot}(A_{*,*})$ with
$\mathrm{Tot}(A_{*,*})_{n}=\bigoplus_{p+q=n}A_{p,q}$
where the boundary maps given by
$\partial_{n}(x)=\partial^{h}_{p}(x)+(-1)^{p}\partial_{q}^{v}(x)$
for every $x\in X_{p,q}$ with $p+q=n$.
Another way to get a chain complex from a bisimplicial $R$-module AA is to consider the simplicial $R$-module $\mathrm{diag}A$ and take its Moore complex $(\mathrm{diag}A)_{*}$. By the Dold-Puppe theorem, these two different constructions give chain complexes that are homotopic.
Theorem 4.3 (Dold-Puppe Theorem [4]).
Let $A:\Delta^{op}\times\Delta^{op}\to R$-Mod be a bisimplicial $R$-module. Then the inclusion map $(\mathrm{diag}A)_{*}\to\mathrm{Tot}(A_{*,*})$ is a chain homotopy equivalence. This equivalence is natural with respect to morphisms of bisimplicial $R$-modules.
Proof.
See [9, Chp IV, Thm 2.4] for a proof. ∎
The homology of the total complex $\mathrm{Tot}(A_{*,*})$ can be calculated by spectral sequences associated to the double complex. Depending on the chosen filtration, horizontal or vertical, there are two spectral sequences that converge to the homology of the total complex (see [19]):
$\begin{split}{}^{I}E^{2}_{p,q}&=H^{h}_{p}(H^{v}_{q}(A_{*,*}))\Rightarrow H_{p+q}(\mathrm{Tot}(A_{*,*}))\\ {}^{II}E^{2}_{p,q}&=H^{v}_{p}(H^{h}_{q}(A_{*,*}))\Rightarrow H_{p+q}(\mathrm{Tot}(A_{*,*}))\end{split}$ (2)
These spectral sequences are natural with respect to morphisms between double complexes. As a consequence of this naturality, and by applying the Dold-Puppe theorem, we obtain the following.
Proposition 4.4.
Let $f:A\to B$ be a morphism of bisimplicial $R$-modules.
  1. (1)
    If for every $q\geq 0$, the induced map $f_{*}:H^{h}_{*}(A_{*,q})\to H^{h}(B_{*,q})$ is an isomorphism, then $(\mathrm{diag}f)_{*}:H_{*}(\mathrm{diag}A)\to H_{*}(\mathrm{diag}B)$ is an isomorphism.
  2. (2)
    If for every $p\geq 0$, the induced map $f_{*}:H^{v}_{*}(A_{p,*})\to H^{v}(B_{p,*})$ is an isomorphism, then $(\mathrm{diag}f)_{*}:H_{*}(\mathrm{diag}A)\to H_{*}(\mathrm{diag}B)$ is an isomorphism.
Proof.
This follows from the spectral sequences above and from Theorem 4.3. ∎
5. Proof of Theorem 1.2
To prove Theorem 1.2, we will apply the argument used by Cegarra in [3]. We first recall the terminology used in the statement of Theorem 1.2.
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor between two small categories, and $R$ be a commutative ring. We denote by $G:=(\varphi/-):\mathcal{D}\to Cat$ the functor that sends $d\in\mathcal{D}$ to the comma category $\varphi/d$. Let $NG:\mathcal{D}\to sSet$ denote the functor that sends $d\in\mathcal{D}$ to the nerve $NG(d)$. Consider the simplicial replacement $N(\mathcal{D};NG)$ of $NG$. For each $d\in\mathcal{D}$, we can identify $NG(d)_{q}$ with the set of pairs $(\tau,\mu)$ where $\tau=(c_{0}\to\cdots\to c_{q})\in N\mathcal{C}_{q}$ and $\mu:\varphi(c_{q})\to d$ is a morphism in $\mathcal{D}$. This identification allows us to write the $(p,q)$-simplices in $N(\mathcal{D};NG)$ as the triples $(\sigma,\tau,\mu)$ where
$\sigma=(d_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{p}}\,}d_{p})\in N\mathcal{D}_{p},\quad\tau=(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\beta_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\beta_{q}}\,}c_{q})\in N\mathcal{C}_{q}$
and $\mu:\varphi(c_{q})\to d_{0}$ is a morphism in $\mathcal{D}$. If $f_{h}\times f_{v}:([r],[s])\to([p],[q])$ is a morphism in $\Delta\times\Delta$, then the induced map $(f_{h}\times f_{v})^{*}$ is defined by
$(f_{h}\times f_{v})^{*}(\sigma,\tau,\mu)=(f_{h}^{*}(\sigma),f_{v}^{*}(\tau),\mu^{\prime})$ (3)
where $\mu^{\prime}:\varphi(c_{f_{v}(q)})\to d_{f_{h}(0)}$ is the morphism in $\mathcal{D}$ defined as the composition
$\varphi(c_{f_{v}(q)})\xrightarrow{\varphi(\beta_{f_{v}(q)+1})}\cdots\xrightarrow{\varphi(\beta_{q})}\varphi(c_{q})\smash{\,\mathop{\longrightarrow}\limits^{\mu}\,}d_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\xrightarrow{\alpha_{f_{h}(0)}}d_{f_{h}(0)}.$ (4)
This description of induced maps $(f_{h}\times f_{v})^{*}$ follows from the identification of $NG(d_{0})_{q}$ with the set of pairs $(\tau,\mu:\varphi(c_{q})\to d_{0})$, so it respects the composition of functions in $\Delta\times\Delta$.
Lemma 5.1.
Consider $N\mathcal{C}$ as a bisimplicial set by taking $(N\mathcal{C})_{p,q}=N\mathcal{C}_{q}$ for all $p$, and by taking all the horizontal maps as the identity maps. Then the map $\kappa:N(\mathcal{D};NG)\to N\mathcal{C}$ defined by $\kappa(\sigma,\tau,\mu)=\tau$ is a map of bisimplicial sets.
Proof.
This follows from the formula
$(f_{h}\times f_{v})^{*}(\sigma,\tau,\mu)=(f_{h}^{*}(\sigma),f_{v}^{*}(\tau),\mu^{\prime})$
given above in (3). ∎
Let $X:=\operatorname*{hocolim}_{\mathcal{D}}NG$ denote the homotopy colimit of the functor $NG$, and let $\kappa:X\to N\mathcal{C}$ be the map defined by $\kappa(\sigma,\tau,\mu)=\tau$. By Lemma 5.1, $\kappa$ is a simplicial map obtained as a restriction of a bisimplicial map. Our main aim in this section is to prove Theorem 1.2 which states that for any coefficient system $\mathcal{M}:\Delta(N\mathcal{C})\to R$-Mod, the induced map
$\kappa^{*}:H^{*}_{Th}(\mathcal{C};\mathcal{M})\to H^{*}(X;\kappa^{*}\mathcal{M})$
is an isomorphism. For this, we introduce a double complex of projective $R\Delta(N\mathcal{C})$-modules.
Definition 5.2.
For $p,q\geq 0$, let $\mathbb{P}$ denote the bisimplicial $R\Delta(N\mathcal{C})$-module defined by
$\mathbb{P}_{p,q}=\bigoplus_{(\sigma,\tau,\mu)\in N(\mathcal{D};G)_{p,q}}R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?).$
For $\theta\in N\mathcal{C}_{n}$, $\mathbb{P}_{p,q}(\theta)$ is the free $R$-module with basis given by the set
$X_{p,q}(\theta)=\{(\sigma,\tau,\mu,f)\,|\,(\sigma,\tau,\mu)\in N(\mathcal{D};NG)_{p.q},f:[q]\to[n],f^{*}(\theta)=\tau\}.$
The vertical and horizontal maps for $\mathbb{P}$ are defined as follows: If $f_{h}\times f_{v}:([r],[s])\to([p],[q])$ is a morphism in $\Delta\times\Delta$, then for $(\sigma,\tau,\mu,f)\in X_{p,q}(\theta)$, we have
$(f_{h}\times f_{v})^{*}(\sigma,\tau,\mu,f)=(f_{h}^{*}(\sigma),f_{v}^{*}(\tau),\mu^{\prime},f\circ f_{v})$
where $\mu^{\prime}:\varphi(c_{f_{v}(q)})\to d_{f_{h}(0)}$ is the morphism in $\mathcal{D}$ defined in (4).
Proposition 5.3.
The chain complex $(\mathrm{diag}\mathbb{P})_{*}$ defines a projective resolution of $\underline{R}$ as an $R\Delta(N\mathcal{C})$-module.
Proof.
Since $\mathbb{P}$ consists of projective $R\Delta(N\mathcal{C})$-modules, we only need to show that the sequence
$\cdots\to(\mathrm{diag}\mathbb{P})_{k}\to(\mathrm{diag}\mathbb{P})_{k-1}\to\cdots\to(\mathrm{diag}\mathbb{P})_{0}\to\underline{R}\to 0$ (5)
is exact. We will do this by relating this chain complex to another complex. Let $Q$ be the simplicial $R\Delta(N\mathcal{C})$-module such that for each $q\geq 0$,
$Q_{q}=\bigoplus_{\tau\in N\mathcal{C}_{q}}R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?).$
For every simplex $\theta\in N\mathcal{C}_{n}$, $Q_{q}(\theta)$ is a free $R$-module with basis given by
$Y_{q}(\theta)=\{(\tau,f)\,|\,\tau\in N\mathcal{C}_{q},f:[q]\to[n]\text{ such that }f^{*}(\theta)=\tau\}.$
For every morphism $g:[q^{\prime}]\to[q]$ in $\Delta$, the induced map $g^{*}(\theta):Q_{q}(\theta)\to Q_{q^{\prime}}(\theta)$ is defined by
$g^{*}(\theta)(\tau,f)=(g^{*}\tau,f\circ g).$
Let $Q_{*}$ be the Moore complex of the simplicial $R\Delta(N\mathcal{C})$-module $Q$. By the argument used in the proof of Proposition 3.3, the sequence
$\widetilde{Q}_{*}:\cdots\to Q_{k}\smash{\,\mathop{\longrightarrow}\limits^{\partial_{k}}\,}Q_{k-1}\to\cdots\to Q_{0}\smash{\,\mathop{\longrightarrow}\limits^{\epsilon}\,}\underline{R}\to 0$ (6)
is exact. Let $\mathbb{Q}$ be the bisimplicial $R\Delta(N\mathcal{C})$-module such that $\mathbb{Q}_{p,q}=Q_{q}$ for all $p,q\geq 0$, with vertical boundary and degeneracy maps given by the boundary and degeneracy maps of $Q_{*}$, and the horizontal boundary and degeneracy maps given as the identity maps. Consider the bisimplicial map $\xi:\mathbb{P}\to\mathbb{Q}$ defined by $\xi_{p,q}(\theta)(\sigma,\tau,\mu,f)=(\tau,f)$ for every $(\sigma,\tau,\mu,f)\in X_{p,q}(\theta)$.
We claim that for a fixed $q$, the simplicial map $\xi_{*,q}(\theta):\mathbb{P}_{*,q}(\theta)\to\mathbb{Q}_{*,q}(\theta)$ induces a homology isomorphism for every $\theta\in\Delta(X)$. Since the simplicial $R$-module $\mathbb{Q}_{*,q}(\theta)$ is the direct sum of simplicial points indexed by the pairs $(\tau,f)$, it is enough to fix a pair $(\tau,f)$, and show that the simplicial $R$-module $\xi_{*,q}(\theta)^{-1}(\tau,f)$ has the homology of a point.
The simplicial $R$-module $\xi_{*,q}(\theta)^{-1}(\tau,f)$ is isomorphic to the linearization of the simplicial set that is isomorphic to the nerve of the comma category $\mathcal{D}/\varphi(c_{0})$. Since this category has a terminal object, the homology of its nerve has the homology of a point. This proves that for every $q\geq 0$, the simplicial map $\xi_{*,q}(\theta):\mathbb{P}_{*,q}(\theta)\to\mathbb{Q}_{*,q}(\theta)$ induces a homology isomorphism.
[…]
Proof of Theorem 1.2.
By Proposition 5.3,
$(\mathrm{diag}\mathbb{P})_{*}:\cdots\to(\mathrm{diag}\mathbb{P})_{k}\to(\mathrm{diag}\mathbb{P})_{k-1}\to\cdots\to(\mathrm{diag}\mathbb{P})_{0}\to\underline{R}\to 0$
is a projective resolution of $\underline{R}$ as an $R\Delta(N\mathcal{C})$-module. Hence for any functor $\mathcal{M}:\Delta(N\mathcal{C})\to R$-mod, the cohomology of the cochain complex $\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{P})_{*},\mathcal{M})$ is isomorphic to the Thomason cohomology $H^{*}_{Th}(\mathcal{C};\mathcal{M})$. Note that for each $k\geq 0$,
$\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{P})_{k},\mathcal{M})$ $\cong\prod_{(\sigma,\tau,\mu)\in N(\mathcal{D};NG)_{k,k}}\mathrm{Hom}_{R\Delta(N\mathcal{C})}(R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?),\mathcal{M})$
$\cong\prod_{(\sigma,\tau,\mu)\in N(\mathcal{D};NG)_{k,k}}\mathcal{M}(\tau)$
$\cong C^{k}(\mathrm{diag}N(\mathcal{D};NG);\kappa^{*}(\mathcal{M})).$
Let $\mathbb{Q}$ be the bisimplicial $R\Delta(N\mathcal{C})$-module introduced in the proof of Proposition 5.3, and let $\xi:\mathbb{P}\to\mathbb{Q}$ be the bisimplicial map defined by $\xi_{p,q}(\theta)(\sigma,\tau,\mu,f)=(\tau,f)$ for every $(\sigma,\tau,\mu,f)\in X_{p,q}(\theta)$. There is an isomorphism for $\mathrm{diag}\mathbb{Q}$ similar to the above isomorphism:
$\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{Q})_{k},\mathcal{M})$ $\cong\prod_{\tau\in N\mathcal{C}_{k}}\mathrm{Hom}_{R\Delta(N\mathcal{C})}(R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?),\mathcal{M})$
$\cong\prod_{\tau\in N\mathcal{C}_{k}}\mathcal{M}(\tau)\cong C^{k}(N\mathcal{C};\mathcal{M}).$
These two isomorphisms fit into a commuting diagram
${\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{P})_{*},\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${C^{*}(\mathrm{diag}N(\mathcal{D};NG);\kappa^{*}(\mathcal{M}))}$${\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{Q})_{*},\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\xi^{*}}$${C^{*}(N\mathcal{C};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa^{*}}$
where the first vertical map is induced by the bisimplicial map $\xi:\mathbb{P}\to\mathbb{Q}$ and the second vertical map is induced by the simplicial map $\kappa:N(\mathcal{D};NG)\to N\mathcal{C}$. Since the augmented complexes for $(\mathrm{diag}\mathbb{P})_{*}$ and $(\mathrm{diag}\mathbb{Q})_{*}$ are projective resolutions for the constant functor $\underline{R}$, the first vertical map $\xi^{*}$ induces an isomorphism on cohomology. Hence $\kappa^{*}$ also induces an isomorphism on cohomology. ∎
By reversing the direction of morphisms and changing the statements accordingly, it is easy to see that the following version of Theorem 1.2 also holds.
Theorem 5.4.
Let $\varphi:\mathcal{C}\rightarrow\mathcal{D}$ be a functor between two small categories and $\backslash\varphi:\mathcal{D}^{op}\rightarrow Cat$ denote the functor that sends each $d\in\mathcal{D}$ to the comma category $d\backslash\varphi$. Then for every coefficient system $\mathcal{M}:\Delta(N\mathcal{C})\rightarrow R$-Mod, the homotopy equivalence $\upsilon:Y=\operatorname*{hocolim}_{\mathcal{D}^{op}}N(-\backslash\varphi)\to N\mathcal{C}$ induces an isomorphism
$H^{*}_{Th}(\mathcal{C};\mathcal{M})\cong H^{*}(Y;\upsilon^{*}\mathcal{M})$
where $\upsilon^{*}M$ denotes the coefficient system for $Y$ induced by $\upsilon$.
6. Cohomology of bisimplicial sets with nontrivial coefficients
Let $X:\Delta^{op}\times\Delta^{op}\to Sets$ be a bisimplicial set and $R$ a commutative ring. The simplex category of $X$ is the category $\Delta(X)$ whose objects are the simplices $\sigma\in X_{p,q}$ and where a morphism $g\in\mathrm{Mor}_{\Delta(X)}(\tau,\sigma)$ is given by a morphism $g^{\prime}:[r]\times[s]\to[p]\times[q]$ in $\Delta\times\Delta$ such that $(g^{\prime})^{*}(\sigma)=\tau$. For a bisimplicial set $X$, a functor $\mathcal{M}:\Delta(X)\to R$-Mod is called a coefficient system for $X$ over $R$.
The cohomology of a bisimplicial set $X$ with coefficients in a coefficient system $\mathcal{M}$ is defined as the cohomology of the simplex category $\Delta(X)$ with coefficients in $R\Delta(X)$-module $\mathcal{M}$. As we did for the cohomology of simplicial sets, we can define a projective resolution for $\underline{R}$ as an $R\Delta(X)$-module as follows:
Definition 6.1.
Let $\mathbb{P}$ be the bisimplicial $R\Delta(X)$-module such that for every $r,s\geq 0$,
$\mathbb{P}_{r,s}=\bigoplus_{\tau\in X_{r,s}}R\mathrm{Mor}_{\Delta(X)}(\tau,?).$
For every $\theta\in X_{p,q}$, $\mathbb{P}_{r,s}(\theta)$ is a free $R$-module with basis given by the pairs $(\tau,g)$ where $\tau\in X_{r,s}$ and $g\in\mathrm{Mor}_{\Delta(X)}(\tau,\theta)$. For every morphism $f:[r^{\prime}]\times[s^{\prime}]\to[r]\times[s]$ in $\Delta\times\Delta$, $f^{*}:\mathbb{P}_{r,s}(\theta)\to\mathbb{P}_{r^{\prime},s^{\prime}}(\theta)$ is defined by
$f^{*}(\tau,g)=(f^{*}(\tau),g\circ f)$
for every $\tau\in X_{r,s}$ and $g\in\mathrm{Mor}_{\Delta(X)}(\tau,\sigma)$.
Let $P_{*,*}$ denote the double complex of $R\Delta(X)$-modules corresponding to the bisimplicial $R\Delta(X)$-module $\mathbb{P}$.
Lemma 6.2.
The chain complex $\mathrm{Tot}(P_{*,*})\to\underline{R}$ is a projective resolution of $\underline{R}$ as an $R\Delta(X)$-module.
Proof.
Let $\theta\in X_{p,q}$. By Theorem 4.3, the inclusion map $(\mathrm{diag}\mathbb{P}(\theta))_{*}\to\mathrm{Tot}(P_{*,*}(\theta))$ is a chain homotopy equivalence where $(\mathrm{diag}\mathbb{P}(\theta))_{*}$ is the chain complex associated to the simplicial $R$-module $\mathrm{diag}\mathbb{P}(\theta)$. The chain complex $(\mathrm{diag}P(\theta))_{*}$ is isomorphic to the simplicial chain complex of the simplicial set $\Delta[p]\times\Delta[q]$. Since the realization of this simplicial set is homeomorphic to $|\Delta[p]|\times|\Delta[q]|$ which is contractible, we can conclude that $(\mathrm{diag}P(\theta))_{*}\to R$ is exact. This gives that the chain complex $\mathrm{Tot}(P_{*,*}(\theta))\to R$ is exact. Since this is true for all $\theta\in X_{p,q}$, the chain complex $\mathrm{Tot}(P_{*,*})\to\underline{R}$ of $R\Delta(X)$-modules is a projective resolution of $\underline{R}$. ∎
Lemma 6.2 shows that there is a standard cochain complex to calculate the cohomology groups $H^{*}(X;\mathcal{M})$ for a bisimplicial set $X$ with coefficient system $\mathcal{M}$. For each $n\geq 0$, this cochain complex has $n$-cochains
$C^{n}(X;\mathcal{M})=\mathrm{Hom}_{R\Delta(X)}(\mathrm{Tot}(P_{*,*})_{n};\mathcal{M})\cong\prod_{r+s=n}\prod_{\tau\in X_{r,s}}\mathcal{M}(\tau)$
with coboundary maps induced by the simplicial maps of $\mathbb{P}$.
Proposition 6.3.
Let $X$ be a bisimplicial set and $J:\Delta(\mathrm{diag}X)\to\Delta(X)$ be the functor defined by the inclusion of simplices in $\mathrm{diag}X$. For every coefficient system $\mathcal{M}$ for $X$, the induced homomorphism
$J^{*}:H^{*}(X;\mathcal{M})\to H^{*}(\mathrm{diag}X;J^{*}\mathcal{M})$
is an isomorphism.
Proof.
By Lemma 6.2, $\mathrm{Tot}(P_{*,*})\to\underline{R}$ is a projective resolution of $\underline{R}$ as an $R\Delta(X)$-module. By Theorem 4.3, we see that $(\mathrm{diag}\mathbb{P})_{*}\to\underline{R}$ is a also a projective resolution of $\underline{R}$ as an $R\Delta(X)$-module, hence the chain map defined by the inclusion of diagonal modules induces a chain homotopy equivalence $j_{*}:(\mathrm{diag}\mathbb{P})_{*}\to\mathrm{Tot}(\mathbb{P}_{*,*})$. We have
$\mathrm{Hom}_{\Delta(X)}((\mathrm{diag}\mathbb{P})_{n},\mathcal{M})\cong\prod_{x\in X_{n,n}}\mathcal{M}(x)\cong C^{n}(\mathrm{diag}X;J^{*}\mathcal{M})$
where $C^{n}(\mathrm{diag}X;J^{*}\mathcal{M})$ is the cochain module of the simplicial set $\mathrm{diag}X$. The coboundary maps are compatible, so we have an isomorphism of cochain complexes
$C^{*}(\mathrm{diag}X;J^{*}\mathcal{M})\cong\mathrm{Hom}_{\Delta(X)}((\mathrm{diag}\mathbb{P})_{*},\mathcal{M})).$
Applying the hom-functor to $j^{*}$, we obtain a commuting diagram of cochain complexes
[…]
where the horizontal maps are the chain isomorphisms described above, and the second vertical map is a chain homotopy equivalence. This means $J^{*}$ is also a chain homotopy equivalence, hence $J^{*}$ induces isomorphism on cohomology. ∎
The following definition will be important for our applications with bisimplicial sets.
Definition 6.4.
Let $X$ be a bisimplicial set. A coefficient system on $X$ is called horizontally constant if for every $f:[p^{\prime}]\to[p]$ and every $x\in X_{p,q}$, the induced map
$\mathcal{M}(f\times\mathrm{id}):\mathcal{M}(x)\to\mathcal{M}((f\times\mathrm{id})^{*}(x))$
is the identity map. A coefficient system $\mathcal{M}$ on $\mathrm{diag}X$ is called strongly extendable if there is a horizontally constant coefficient system $\mathcal{M}^{\prime}$ on $\Delta(X)$ such that $\mathcal{M}=J^{*}(\mathcal{M}^{\prime})$.
We have the following observation.
Lemma 6.5.
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor and $\kappa:X=\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\to N\mathcal{C}$ be the simplicial map defined in Definition 1.1. Then for every coefficient system $\mathcal{M}$ for $N\mathcal{C}$, the induced system $\kappa^{*}\mathcal{M}$ is a strongly extendable coefficient system for $X$.
Proof.
By Lemma 5.1, the simplicial map $\kappa$ is the restriction of a bisimplicial map defined by $\kappa(\sigma,\tau,\mu)=\tau$. For every $f:[p^{\prime}]\to[p]$, we have
$(\kappa^{*}\mathcal{M})(f\times\mathrm{id})=\mathcal{M}(\kappa((f\times\mathrm{id})^{*}))=\mathcal{M}(\mathrm{id})=\mathrm{id}$
by Equation 3. This shows that the bisimplicial module $\kappa^{*}M$ defined on $N(\mathcal{D},N(\varphi/-))$ is horizontally constant. Hence $\kappa^{*}\mathcal{M}$ is strongly extendable. ∎
Let $F:\mathcal{D}\to sSet$ be a functor. The homotopy colimit $\operatorname*{hocolim}_{\mathcal{D}}F$ is the diagonal of the bisimplicial set $N(\mathcal{D};F)$ whose $(p,q)$-simplices are given by pairs $(\sigma,\tau)$ where $\sigma=(d_{0}\to\cdots\to d_{p})\in N\mathcal{D}_{p}$ and $\tau\in F(d_{0})_{q}$. A coefficient system $\mathcal{M}$ for $\operatorname*{hocolim}_{\mathcal{D}}F$ is a functor from the simplex category $\Delta(\operatorname*{hocolim}_{\mathcal{D}}F)$ to the category of $R$-modules.
Lemma 6.6.
Let $U:F_{1}\to F_{2}$ be a natural transformation between functors $F_{1},F_{2}:\mathcal{D}\to Cat$. Then $U$ induces a bisimplicial map
$\mathfrak{U}:N(\mathcal{D};NF_{1})\to N(\mathcal{D};NF_{2}),$
hence a simplicial map $\mathfrak{u}:\operatorname*{hocolim}_{\mathcal{D}}NF_{1}\to\operatorname*{hocolim}_{\mathcal{D}}NF_{2}$.
Proof.
Let $\mathfrak{U}$ be the map that sends a pair $(\sigma,\tau)$ where
$\sigma=(d_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{p}}\,}d_{p})\in N\mathcal{D}_{p}\text{ and }\tau=(x_{0}\smash{\,\mathop{\longrightarrow}\limits^{\beta_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\beta_{q}}\,}x_{q})\in NF_{1}(d_{0})_{q},$
to the pair $(\sigma,\tau^{\prime})$ where
$\tau^{\prime}=U(d_{0})(\tau):=(U(d_{0})(x_{0})\xrightarrow{U(d_{0})(\beta_{1})}U(d_{0})(x_{1})\xrightarrow{}\cdots\xrightarrow{U(d_{0})(\beta_{q})}U(d_{0})(x_{q}))$
in $N(F_{2}(d_{0}))_{q}$. For every morphism $f_{h}\times f_{v}:([p^{\prime}],[q^{\prime}])\to([p],[q])$ in $\Delta\times\Delta$, and for $(\sigma,\tau)\in N(\mathcal{D};NF_{1})$, we have
$\bigl((f_{h}\times f_{v})^{*}\circ\mathfrak{U}\bigr)(\sigma,\tau)=(f_{h}^{*}(\sigma),F_{2}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau^{\prime}))$
and
$\bigl(\mathfrak{U}\circ(f_{h}\times f_{v})^{*}\bigr)(\sigma,\tau)=(f_{h}^{*}(\sigma),\bigl(F_{1}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))\bigr)^{\prime}).$
Note that
$F_{2}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau^{\prime}))$ $=F_{2}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(U(d_{0})(\tau)))$
$=F_{2}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(U(d_{0})(f_{v}^{*}(\tau)))$
$=U(d_{f_{h}(0)})\bigl(F_{1}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))\bigr)$
$=\bigl(F_{1}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))\bigr)^{\prime}$
where the second equality holds because $U(d_{0}):F_{1}(d_{0})\to F_{2}(d_{0})$ is a functor, and the third equality holds because $U$ is a natural transformation. So the equality
$(f_{h}\times f_{v})^{*}\circ\mathfrak{U}=\mathfrak{U}\circ(f_{h}\times f_{v})^{*}$
holds. ∎
For every $d\in\mathcal{D}$, there is a simplicial map $i_{d}:NF(d)\to\operatorname*{hocolim}_{\mathcal{D}}F$ that sends a simplex $\tau\in NF(d)_{p}$ to the pair $((d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d),\tau)$ in $N(\mathcal{D};NF)_{p,p}$.
Proposition 6.7.
Let $U:F_{1}\to F_{2}$ be a natural transformation between functors $F_{1},F_{2}:\mathcal{D}\to Cat$, and let $\mathcal{M}$ be a strongly extendable coefficient system defined on $\operatorname*{hocolim}_{\mathcal{D}}F_{2}$. If for every $d\in\mathcal{D}$, the induced map
$NU(d)^{*}:H^{*}(NF_{2}(d),i_{d}^{*}\mathcal{M})\to H^{*}(NF_{1}(d);NU(d)^{*}i_{d}^{*}\mathcal{M})$
is an isomorphism, then
$\mathfrak{u}^{*}:H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{2};\mathcal{M})\to H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{1};\mathfrak{u}^{*}\mathcal{M})$
is an isomorphism.
Proof.
Let $\mathcal{M}^{\prime}$ be a horizontally constant coefficient system for the bisimplicial set $N(\mathcal{D};NF_{2})$ such that $J^{*}(\mathcal{M}^{\prime})=\mathcal{M}$. By Lemma 6.6, the natural transformation $U:F_{1}\to F_{2}$ induces a bisimplicial map $\mathfrak{U}:N(\mathcal{D};NF_{1})\to N(\mathcal{D};NF_{2})$. Consider the following commuting diagram:
${H^{*}(N(\mathcal{D};NF_{2});\mathcal{M}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{U}^{*}}$$\scriptstyle{J^{*}}$${H^{*}(N(\mathcal{D};NF_{1});\mathfrak{U}^{*}\mathcal{M}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J^{*}}$${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{2};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{u}^{*}}$${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{1};\mathfrak{u}^{*}\mathcal{M})}$
where horizontal maps are defined as in Lemma 3.4 and the vertical maps are the maps defined in Proposition 6.3. By Proposition 6.3, vertical maps $J^{*}$ are isomorphisms. Hence $\mathfrak{u}^{*}$ is an isomorphism if $\mathfrak{U}^{*}$ is an isomorphism.
Let $X_{i}=N(\mathcal{D};F_{i})$ for $i=1,2$. The homomorphism $\mathfrak{U}^{*}$ is defined on the cochain level as the chain map $\mathfrak{U}^{*}:C^{*}(X_{2};\mathcal{M}^{\prime})\to C^{*}(X_{1};\mathfrak{U}^{*}\mathcal{M}^{\prime})$ which sends $f\in C^{n}(X_{2};\mathcal{M}^{\prime})$ to $\mathfrak{U}^{*}(f)\in C^{n}(X_{1};\mathfrak{U}^{*}\mathcal{M}^{\prime})$ such that
$(\mathfrak{U}^{*}f)(\sigma,\tau)=f(\sigma,NU(d_{0})(\tau))$
for every simplex $(\sigma,\tau)$ in $X_{1}=N(\mathcal{D};NF_{1})$.
Recall that for a bisimplicial set $X$ and a coefficient system $\mathcal{M}$, the cochain complex $C^{*}(X;\mathcal{M})$ has $n$-cochains
$C^{n}(X;\mathcal{M})=\mathrm{Hom}_{R\Delta(X)}(\mathrm{Tot}(P_{*,*})_{n};\mathcal{M})\cong\prod_{p+q=n}\prod_{x\in X_{p,q}}\mathcal{M}(x)$
with coboundary maps induced by the simplicial maps of $\mathbb{P}$. We can consider $C^{*}(X;\mathcal{M})$ as the total complex of the bi-cosimplicial $R$-module $\mathbb{C}$ with $\mathbb{C}^{p,q}(X;\mathcal{M})\cong\prod_{x\in X_{p,q}}\mathcal{M}(x)$.
The cohomology of the associated total complex $\mathrm{Tot}(C^{*,*})$ can be calculated by spectral sequences associated to the double complex as discussed earlier in Section 4. There are two spectral sequences that converge to the cohomology of the total complex (see [19]):
$\begin{split}{}^{I}E_{2}^{p,q}&=H_{h}^{p}(H_{v}^{q}(C^{*,*}))\Rightarrow H^{p+q}(\mathrm{Tot}(C^{*,*}))\\ {}^{II}E_{2}^{p,q}&=H_{v}^{p}(H_{h}^{q}(C^{*,*}))\Rightarrow H^{p+q}(\mathrm{Tot}(C^{*,*})).\end{split}$ (7)
These spectral sequences are natural with respect to morphisms between double complexes. Hence $\mathfrak{U}^{*}$ induces a morphism of spectral sequences
$\mathfrak{U}^{*}:{}^{I}E_{2}^{p,q}(X_{2};\mathcal{M}^{\prime})\to{}^{I}E_{2}^{p,q}(X_{1};\mathfrak{U}^{*}\mathcal{M}^{\prime})$
which will be an isomorphism if for every fixed $p$, the induced map
$\mathfrak{U}^{*}:H_{v}^{*}(X^{p,*}_{2};\mathcal{M}^{\prime})\to H_{v}^{*}(X^{p,*}_{1};\mathfrak{U}^{*}\mathcal{M}^{\prime})$
is an isomorphism. The simplicial set $X_{2}^{p,*}$ is isomorphic to the disjoint union $\coprod_{\sigma\in N\mathcal{D}_{p}}NF_{2}(d_{0})$ by Equation 1. Hence $H^{*}_{v}(X_{2}^{p,*};\mathcal{M}^{\prime})$ is isomorphic to the direct product
$\prod_{\sigma\in N\mathcal{D}_{p}}H^{*}(NF_{2}(d_{0}),i_{\sigma}^{*}\mathcal{M}^{\prime})$
where $i_{\sigma}:NF_{2}(d_{0})\to X_{2}^{p,*}$ is the simplicial map that sends $\tau$ to $(\sigma,\tau).$ Since $\mathcal{M}^{\prime}$ is horizontally constant, $i_{\sigma}^{*}\mathcal{M}^{\prime}=i_{d_{0}}^{*}\mathcal{M}$ as coefficients systems for $NF_{2}(d_{0})$. Hence $\mathfrak{U}^{*}$ is a disjoint union of homomorphisms
$NU(d_{0})^{*}:H^{*}(NF_{2}(d_{0});i_{d_{0}}^{*}\mathcal{M})\to H^{*}(NF_{1}(d_{0});NU(d_{0})^{*}i_{d_{0}}^{*}\mathcal{M})$
over all $\sigma:(d_{0}\to\cdots\to d_{p})$ in $N\mathcal{D}_{p}$. Each of these homomorphisms are isomorphism by the assumption of the proposition. This gives that $\mathfrak{U}^{*}$ is an isomorphism, hence $\mathfrak{u}^{*}$ is also an isomorphism. ∎
As a consequence of Theorem 1.2, we will prove a cohomological version of Quillen’s Theorem A stated as Theorem 1.3 in the introduction.
Proof of Theorem 1.3.
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor between small categories and $\mathcal{M}$ be any coefficient system for $N\mathcal{D}$. Consider the natural transformation
$U:\varphi/-\to\mathrm{id}/-,$
defined such that for each $d\in\mathcal{D}$, $U(d):\varphi/d\to\mathrm{id}/d$ is the functor that sends the pair $(c,\mu)\in\varphi/d$ to $(\varphi(c),\mu)\in\mathrm{id}/d$. For each $d\in\mathcal{D}$, there is a commuting diagram of simplicial maps
${N(\varphi/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}}$$\scriptstyle{NU(d)}$${N\mathcal{C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{N\varphi}$${N(\mathrm{id}/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}^{\prime}}$${N\mathcal{D}}$
where $j_{d}$ and $j_{d}^{\prime}$ are the simplicial maps defined in Definition 1.1.
For each $d\in\mathcal{D}$, let $\mathcal{M}_{d}=(j_{d}^{\prime})^{*}\mathcal{M}$ and $\mathcal{M}_{d}^{\prime}=(NU(d))^{*}\mathcal{M}_{d}$. We assume that for every $d\in\mathcal{D}$, the simplicial map $NU(d):N(\varphi/d)\to N(\mathrm{id}/d)$ induces an isomorphism
$H^{*}(N(\mathrm{id}/d);\mathcal{M}_{d})\cong H^{*}(N(\varphi/d);\mathcal{M}_{d}^{\prime}).$
Let $\kappa:\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\to N\mathcal{C}$ and $\kappa^{\prime}:\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-)\to N\mathcal{D}$ be the simplicial maps defined in Definition 1.1. There is a commuting diagram of simplicial sets
${\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{u}}$$\scriptstyle{\kappa}$${N\mathcal{C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{N\varphi}$${\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa^{\prime}}$${N\mathcal{D}}$
where $\mathfrak{u}$ is the simplicial map induced by the functor $U$ defined above. This diagram induces a commuting diagram of homomorphisms
${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-);\kappa^{*}(N\varphi)^{*}\mathcal{M})}$${H^{*}(N\mathcal{C};(N\varphi)^{*}\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa^{*}}$${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-);(\kappa^{\prime})^{*}\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{u}^{*}}$${H^{*}(N\mathcal{D};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{(\kappa^{\prime})^{*}}$$\scriptstyle{(N\varphi)^{*}}$
By Theorem 1.2, the induced maps $\kappa^{*}$ and $(\kappa^{\prime})^{*}$ are both isomorphisms. We claim that $\mathfrak{u}^{*}$ is also an isomorphism. This will imply that $(N\varphi)^{*}$ is an isomorphism and complete the proof.
To see that $\mathfrak{u}^{*}$ is an isomorphism, we will apply Proposition 6.7. Note that by Lemma 6.5, for any coefficient system $\mathcal{M}$ for $N\mathcal{D}$, the induced coefficient system $(\kappa^{\prime})^{*}\mathcal{M}$ is strongly extendable. Let $i_{d}:N(\varphi/d)\to\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)$ be the simplicial map which sends a simplex $(\tau,\mu)\in N(\varphi/d)_{q}$ to the pair $((d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d),\tau,\mu)$ in $N(\mathcal{D};N(\varphi/-))_{q,q}$, and let $i_{d}^{\prime}:N(\mathrm{id}/d)\to\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-)$ be the simplicial map defined in a similar way for the identity functor. We have
$(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M}=(j_{d}^{\prime})^{*}\mathcal{M}=\mathcal{M}_{d}$
and
$i_{d}^{*}\kappa^{*}(N\varphi)^{*}\mathcal{M}=i_{d}^{*}\mathfrak{u}^{*}(\kappa^{\prime})^{*}\mathcal{M}=NU(d)^{*}(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M}=NU(d)^{*}\mathcal{M}_{d}.$
This gives that for every $d\in\mathcal{D}$,
$NU(d)^{*}:H^{*}(\mathrm{id}/d;(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M})\to H^{*}(\varphi/d;NU(d)^{*}(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M})$
is an isomorphism by the assumption of the theorem. Hence we can apply Proposition 6.7 to the natural transformation $U:\varphi/-\to\mathrm{id}/$-, and conclude that $\mathfrak{u}^{*}$ is an isomorphism. We conclude that $(N\varphi)^{*}$ is an isomorphism. ∎
Quillen’s Theorem A states that if $\varphi:\mathcal{C}\to\mathcal{D}$ is a functor such that for every $d\in\mathcal{D}$, the comma category $\varphi/d$ is contractible, then $N\varphi:N\mathcal{C}\to N\mathcal{D}$ is a homotopy equivalence. When $\varphi/d$ is contractible, the simplicial map $NU(d):N(\varphi/d)\to N(\mathrm{id}/d)$ is a homotopy equivalence, but in general it may not induce an isomorphism on cohomology with arbitrary coefficients. This means that under the hypothesis of Quillen’s Theorem A, in general we cannot conclude that the induced map $\varphi^{*}$ on cohomology is an isomorphism.
The following example due to Husainov [12] shows that under the hypothesis of Quillen’s Theorem A, the induced map $\varphi^{*}:H_{Th}^{*}(\mathcal{D};\mathcal{M})\to H_{Th}^{*}(\mathcal{C};\varphi^{*}\mathcal{M})$ is not an isomorphism for some coefficient systems.
Example 6.8.
Let $\mathcal{C}$ be the category with one object $\{0\}$, and $\mathcal{D}$ be the category with objects $\{0,1\}$ and with one non-identity morphism $\alpha:0\to 1$. Let $\varphi:\mathcal{C}\to\mathcal{D}$ be the inclusion functor. Then for every $d\in\mathcal{D}$, the comma category $\varphi/d$ is contractible. Consider the functor between the factorization categories $\mathfrak{F}\varphi:\mathfrak{F}\mathcal{C}\to\mathfrak{F}\mathcal{D}$. Let $M$ be the natural system over $\mathbb{Z}$ with values $M(\mathrm{id}_{0})=M(\mathrm{id}_{1})=0$ and $M(\alpha)\cong\mathbb{Z}$. It is easy to see that $H^{1}_{BW}(\mathcal{C};M)=0$ and $H^{1}_{BW}(\mathcal{D};M)\cong\mathbb{Z}$, so the homomorphism induced by $\varphi$ on Baues-Wirsching cohomology is not an isomorphism (see [12, Example 5.3] for details). Hence the induced map on Thomason cohomology $\varphi^{*}:H^{*}_{Th}(\mathcal{D};\mathcal{M})\to H^{*}_{Th}(\mathcal{C};\varphi^{*}\mathcal{M})$ is not an isomorphism for all coefficient systems $\mathcal{M}$.
In [12], Husainov gives criteria for a functor between two small categories to induce isomorphisms for different definitions of cohomology of small categories. He also constructs counterexamples to some statements that appear in the literature. Husainov’s examples show that for an adjoint pair $(l,r)$ such that $l:\mathcal{C}\rightleftarrows\mathcal{D}:r$, the induced map
$l^{*}:H^{*}_{Th}(\mathcal{D};\mathcal{M})\to H^{*}_{Th}(\mathcal{C};l^{*}(\mathcal{M}))$
is not always an isomorphism. His examples show in particular that [15, Lem 2.2] and [7, Prop 1.11 and Cor 2.3] do not hold in the way they are stated.
7. Thomason cohomology of the Grothendieck construction
Definition 7.1.
Let $F:\mathcal{D}\to Cat$ be a functor. The Grothendieck construction of $F$ is the category $\int_{\mathcal{D}}F$ whose objects are the pairs $(d,x)$ with $x\in F(d)$, and whose morphisms $(d,x)\to(d^{\prime},x^{\prime})$ are given by pairs $(\alpha,\gamma)$ where $\alpha:d\to d^{\prime}$ is a morphism in $\mathcal{D}$, and $\gamma:F(\alpha)(x)\to x^{\prime}$ is a morphism in $F(d^{\prime})$.
For a functor $F:\mathcal{D}\to Cat$, let $NF:\mathcal{D}\to sSet$ denote the functor that sends $d\in\mathcal{D}$ to the nerve of the category $F(d)$. Consider the bisimplicial set $N(\mathcal{D},NF)$ where for each $p,q\geq 0$,
$N(\mathcal{D};NF)_{p,q}=\coprod_{\sigma=(d_{0}\to\cdots\to d_{p})\in N\mathcal{D}_{p}}NF(d_{0})_{q}.$
The homotopy colimit $\operatorname*{hocolim}_{\mathcal{D}}NF$ is the diagonal simplicial set $\mathrm{diag}N(\mathcal{D};NF)$.
Let $\mathcal{C}=\int_{\mathcal{D}}F$. It is shown by Thomason [18] that there is a simplicial map
$\eta:\operatorname*{hocolim}_{\mathcal{D}}NF\to N\mathcal{C}$
which is a homotopy equivalence. To prove this result, Thomason uses a two step approach. There is a functor $\pi:\mathcal{C}\to\mathcal{D}$ that sends a pair $(d,x)\in\mathcal{C}$ to $d\in\mathcal{D}$. Consider the functor $\widetilde{F}:\mathcal{D}\to Cat$ that sends $d\in\mathcal{D}$ to the comma category $\pi/d$. There are simplicial maps $\lambda_{1}$ and $\lambda_{2}$ that make the following diagram commute up to homotopy:
$\operatorname*{hocolim}\limits_{\mathcal{D}}NF$$\operatorname*{hocolim}\limits_{\mathcal{D}}N\widetilde{F}$$N\mathcal{C}.$$\scriptstyle{\eta}$$\scriptstyle{\lambda_{1}}$$\scriptstyle{\lambda_{2}}$
The simplicial maps $\lambda_{1}$ and $\lambda_{2}$ are defined as the diagonalization of the bisimplicial maps
$N(\mathcal{D};NF)$${N(\mathcal{D};N\widetilde{F})}$$N\mathcal{C},$$\scriptstyle{\lambda^{\prime}_{1}}$$\scriptstyle{\lambda^{\prime}_{2}}$
[…]
$\sigma=(d_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{p}}\,}d_{p})\in N\mathcal{D}_{p},\quad\quad\tau=((d_{0}^{\prime},x_{0})\smash{\,\mathop{\longrightarrow}\limits^{(\beta_{1},\gamma_{1})}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{(\beta_{q},\gamma_{q})}\,}(d_{q}^{\prime},x_{q}))\in N\mathcal{C}_{q},$
and $\mu:d_{q}^{\prime}\to d_{0}$ is a morphism in $\mathcal{D}$.
Definition 7.2.
The bisimplicial map $\lambda^{\prime}_{2}:N(\mathcal{D};N\widetilde{F})\to N\mathcal{C}$ is defined by $\lambda^{\prime}_{2}(\sigma,\tau,\mu)=\tau$ for every $(\sigma,\tau,\mu)\in N(\mathcal{D};N\widetilde{F})_{p,q}$. The bisimplicial map $\lambda^{\prime}_{1}:N(\mathcal{D};N\widetilde{F})\to N(\mathcal{D};NF)$ is induced by the natural transformation $\zeta:\widetilde{F}\to F$, where for each $d\in\mathcal{D}$, the functor $\zeta_{d}:\widetilde{F}(d)=\pi/d\to F(d)$ is defined by $\zeta_{d}\bigl((d^{\prime},x),\mu:d^{\prime}\to d\bigr)=F(\mu)(x)$.
Thomason [18] shows that the simplicial maps $\lambda_{1}$ and $\lambda_{2}$ are both homotopy equivalences. Then he constructs a homotopy between the simplicial maps $\eta\circ\lambda_{1}$, and $\lambda_{2}$, and concludes that $\eta$ is also a homotopy equivalence.
Our main aim in this section is to construct a spectral sequence for the Thomason cohomology of the Grothendieck construction $\int_{\mathcal{D}}F$. We start with the following observation which is an immediate consequence of Theorem 1.2.
Proposition 7.3.
Let $F:\mathcal{D}\to Cat$ be a functor, and let $\mathcal{C}:=\int_{\mathcal{D}}F$ be the Grothendieck construction of $F$. For every coefficient system $\mathcal{M}$ on $N\mathcal{C}$, the homotopy equivalence $\lambda_{2}$ induces an isomorphism
$\lambda_{2}^{*}:H^{*}_{Th}(N\mathcal{C};\mathcal{M})\smash{\,\mathop{\longrightarrow}\limits^{\cong}\,}H^{*}(\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F};\lambda_{2}^{*}\mathcal{M}).$
Proof.
This follows from Theorem 1.2 by applying it to the functor $\pi:\int_{\mathcal{D}}F\to\mathcal{D}$. ∎
To be able to use this proposition in our proofs, we need to show that $\lambda_{2}$ also induces an isomorphism between the cohomology groups of the associated bisimplicial sets. Let $N_{b}\mathcal{C}$ denote the bisimplicial set obtained by considering $N\mathcal{C}$ as a bisimplicial set that is constant in the horizontal direction. Note that every coefficient system $\mathcal{M}$ on $N\mathcal{C}$ is extendable to the bisimplicial set $N_{b}\mathcal{C}$ by taking all the horizontal induced maps to be the identity. We denote the extension of $\mathcal{M}$ also by $\mathcal{M}$. We have the following:
Proposition 7.4.
Let $F:\mathcal{D}\to Cat$ be a functor and $\mathcal{C}=\int_{\mathcal{D}}F$ be the Grothendieck construction of $F$. For every coefficient system $\mathcal{M}$ on $N\mathcal{C}$, the induced map
$(\lambda^{\prime}_{2})^{*}:H^{*}(N_{b}\mathcal{C};\mathcal{M})\smash{\,\mathop{\longrightarrow}\limits\,}H^{*}(N(\mathcal{D};N\widetilde{F});(\lambda^{\prime}_{2})^{*}\mathcal{M})$
is an isomorphism.
Proof.
Consider the following commutative diagram of functors
${\Delta(\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces)}$$\scriptstyle{J}$$\scriptstyle{\lambda_{2}}$${\Delta(N\mathcal{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J}$${\Delta(N(\mathcal{D};N\widetilde{F}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda_{2}^{\prime}}$${\Delta(N_{b}\mathcal{C})}$
where $J$ is the inclusion map of diagonal simplices. This diagram induces the commutative diagram of cohomology groups
${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F};\lambda_{2}^{*}\mathcal{M})}$${H^{*}_{Th}(\mathcal{C};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda_{2}^{*}}$${H^{*}(N(\mathcal{D};N\widetilde{F});(\lambda_{2}^{\prime})^{*}\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J^{*}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces H^{*}(\Delta(N_{b}\mathcal{C});\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{(\lambda_{2}^{\prime})^{*}}$$\scriptstyle{J^{*}}$
By Proposition 7.3 the induced map $\lambda_{2}^{*}$ is an isomorphism, and by Proposition 6.3, both of the maps $J^{*}$ are isomorphisms. Hence $(\lambda_{2}^{\prime})^{*}$ is an isomorphism. ∎
Now we are ready to prove Theorem 1.4.
Proof of Theorem 1.4.
Let $F:\mathcal{D}\to Cat$ be a functor and $\mathcal{C}:=\int_{\mathcal{D}}F$ be the Grothendieck construction of $F$. Let $\pi:\mathcal{C}\to\mathcal{D}$ be the canonical functor that sends a pair $(d,x)\in\int_{\mathcal{D}}F$ to $d\in\mathcal{D}$, and $\widetilde{F}:\mathcal{D}\to Cat$ be the functor that sends $d\in\mathcal{D}$ to the comma category $\pi/d$.
[…]
For a bisimplicial set $X$ and a coefficient system $\mathcal{M}$ for $X$, the cohomology $H^{*}(X;\mathcal{M})$ is defined as the cohomology of the cochain complex $C^{*}(X;\mathcal{M})$ where for each $n\geq 0$,
$C^{n}(X;\mathcal{M})=\mathrm{Hom}_{R\Delta(X)}(\mathrm{Tot}(P_{*,*})_{n};\mathcal{M})\cong\prod_{p+q=n}\prod_{x\in X_{p,q}}\mathcal{M}(x)$
with coboundary maps induced by the simplicial maps of $\mathbb{P}$ introduced in Definition 6.1. The complex $C^{*}(X;\mathcal{M})$ is the total complex of the bi-cosimplicial $R$-module $\mathbb{C}$ with $\mathbb{C}^{p,q}(X;\mathcal{M})\cong\prod_{x\in X_{p,q}}\mathcal{M}(x)$. This allows us to calculate the cohomology of $C^{*}(X;\mathcal{M}_{X})$ using a spectral sequence associated to the double complex. There are two spectral sequences that converge to the cohomology of the total complex of a double complex $C^{*,*}=C^{*,*}(X;\mathcal{M})$ as stated in (7). Here we consider the first one which is of the form:
$E_{2}^{p,q}=H_{h}^{p}(H_{v}^{q}(C^{*,*}))\Rightarrow H^{p+q}(\mathrm{Tot}(C^{*,*})).$
[…]
$\prod_{\sigma\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0}),i_{\sigma}^{*}\mathcal{M}^{\prime}).$
Since $\mathcal{M}^{\prime}$ is horizontally constant, $i_{\sigma}^{*}\mathcal{M}^{\prime}=i_{d_{0}}^{*}J^{*}\mathcal{M}^{\prime}$ where
$i_{d_{0}}:N\widetilde{F}(d_{0})\to\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F}$
[…]
$H^{q}_{v}(C^{p,*})\cong\prod_{\sigma\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0});i_{d_{0}}^{*}J^{*}\mathcal{M}^{\prime}).$
For every $d\in\mathcal{D}$, let $j_{d}:N\widetilde{F}(d)=N(\pi/d)\to N(\mathcal{C})$ be the simplicial map induced by the canonical functor $\pi/d\to\mathcal{C}$ which sends $(c,\mu)$ to $c$. From the definitions we see that the composition
$N\widetilde{F}(d)=N(\pi/d)\smash{\,\mathop{\longrightarrow}\limits^{i_{d}}\,}\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F}\smash{\,\mathop{\longrightarrow}\limits^{J}\,}N(\mathcal{D};N\widetilde{F})\smash{\,\mathop{\longrightarrow}\limits^{\lambda_{2}^{\prime}}\,}N\mathcal{C}$
is equal to $j_{d}$. Hence we can write
$H^{q}_{v}(C^{p,*})\cong\prod_{\sigma\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0});j_{d_{0}}^{*}\mathcal{M}).$
[…]
As a special case of the spectral sequence in Theorem 1.4, consider the following situation: Let $G$ be a group and $X$ be a $G$-simplicial set. In this case the simplex category $\Delta(X)$ is a $G$-category. Let $\mathcal{G}$ denote the one object group category whose morphisms are given by $G$, and let $F:\mathcal{G}\to Cat$ be the functor that sends $\ast\in\mathcal{G}$ to $\Delta(X)$. We denote the Grothendieck construction $\int_{\mathcal{G}}F$ by $\Delta(X)_{G}$.
The objects of $\Delta(X)_{G}$ are the pairs $(\ast,\sigma)$ with $\sigma\in\Delta(X)$. For simplicity we denote the objects of $\Delta(X)_{G}$ as $\sigma$. A morphism $\tau\to\sigma$, where $\sigma\in X_{n}$ and $\tau\in X_{m}$, is given by a pair $(g,f)$ where $g\in G$ and $f:[m]\to[n]$ is a morphism in $\Delta$ such that $f^{*}(\sigma)=g\tau$. A functor $\mathcal{M}:\Delta(X)_{G}\to R$-Mod is called a $G$-equivariant coefficient system on $X$.
If we apply Theorem 1.4 to the category $\Delta(X)_{G}$, we obtain the following corollary.
Corollary 7.5.
Let $X$ be a $G$-simplicial set and $\mathcal{M}$ an equivariant coefficient system on $X$. Then there is a spectral sequence
$E_{2}^{p,q}=H^{p}_{Th}(\mathcal{G};\mathcal{H}^{q}_{\mathcal{M}})\Rightarrow H^{p+q}_{Th}(\Delta(X)_{G};\mathcal{M})$
where $\mathcal{H}^{q}_{\mathcal{M}}$ is the coefficient system on $N\mathcal{G}$ that sends a simplex $\ast\smash{\,\mathop{\longrightarrow}\limits^{g_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{g_{p}}\,}*$ in $N\mathcal{G}$ to the $R$-module $H^{q}(N(\pi/*);(j_{*})^{*}\mathcal{M})$ where $\pi:\Delta(X)_{G}\to\mathcal{G}$ is the canonical projection functor.
Remark 7.6.
As a consequence of Thomason’s homotopy colimit theorem, the nerve of the category $\Delta(X)_{G}$ is homotopy equivalent to
$\operatorname*{hocolim}_{\mathcal{G}}N\Delta(X)$
which is homotopy equivalent to the Borel construction $EG\times_{G}X$ (see [11, Lemma 2.3]). This means that when the coefficient system $\mathcal{M}$ is the constant coefficient system $\underline{R}$, the above spectral sequence becomes the usual Serre spectral sequence
$E_{2}^{p,q}=H^{p}(G;H^{q}(X;R))\Rightarrow H^{p+q}(EG\times_{G}X;R).$
In this sense the above spectral sequence can be considered as the generalization of the Serre spectral sequence to the cohomology of $\Delta(X)_{G}$ with arbitrary $G$-equivariant coefficient systems.
Remark 7.7.
The category $\Delta(X)_{G}$ is often used in the definition of $G$-equivariant cohomology groups $H^{*}_{G}(X;\mathcal{M})$ with coefficients in $\mathcal{M}:\Delta(X)_{G}\to R$-Mod. The $G$-equivariant cohomology groups of $X$ are defined as the cohomology groups of the cochain complex $C^{*}(X;\mathcal{M})$ where, for each $n\geq 0$, the $n$-th chain module is defined by
$C^{n}_{G}(X;\mathcal{M})=\Bigl(\prod_{\sigma\in X_{n}}\mathcal{M}(\sigma)\Bigr)^{G}$
(see [10, Definition 2.3]). Note that in general the equivariant cohomology groups $H^{*}_{G}(X;\mathcal{M})$ are not isomorphic to the Thomason cohomology groups $H^{*}_{Th}(\Delta(X)_{G};\mathcal{M})$ of the category $\Delta(X)_{G}$. In the case where $\mathcal{M}$ is the constant coefficient system $\underline{R}$, the equivariant cohomology $H_{G}^{*}(X;\underline{R})$ is isomorphic to the cohomology $H^{*}(X/G;R)$ of the orbit space $X/G$ whereas the Thomason cohomology $H^{*}_{Th}(\Delta(X)_{G};\underline{R})$ is isomorphic to the cohomology of the Borel construction $H^{*}(EG\times_{G}X;R)$. When the $G$-action on $X$ is not free, these two cohomology groups are not isomorphic in general.
Acknowledgements and Funding Declaration
The first author is supported by TÜBİTAK 2211-A National PhD Scholarship Program. The second author is supported by TÜBİTAK 2219-International Postdoctoral Research Fellowship Program (2023, 2nd term). We thank TÜBİTAK for their support of this research.
We also thank Matthew Gelvin and Caroline Yalçın for reading an earlier version of the paper and for their suggestions. We thank the referee for careful reading of the paper and pointing out a mistake in the earlier version of the paper.
Author contributions statement
All authors contributed to the study conception and design. The first draft of the manuscript was written by both authors and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Data Availability
This manuscript does not report data generation or analysis.
Conflict of interest
The authors declare that they have no conflicts of interest.